2- |x+1|<|2x - 3| Solve the problem with the steps and mark the answers on a chart
1
Expert's answer
2020-12-28T14:43:21-0500
"x+1=0=>x=-1"

"2x-3=0=>x=\dfrac{3}{2}"


If "x<-1:"


"2+x+1<-2x+3=>3x<0=>x<0"

Then "x<-1"


If "-1\leq x<3/2:"


"2-x-1<-2x+3=>x<2"

Then "-1\leq x<3/2"


If "x\geq3/2:"


"2-x-1<2x-3=>3x>4=>x>4\/3"

Then "x>3\/2"

Therefore


"x\in(-\infin, \infin)"